# Single-Channel EEG Tokenization Through Time-Frequency Modeling

### v1

@@ 5.3. Q2: EEG Token Quality Analysis and Frequency Learning
We study the quality of the EEG tokens learned by our TFM-Tokenizer by analyzing four key aspects: (1) token utilization, (2) class-specific distinctiveness, (3) similar class retrieval, and (4) frequency learning capability. We conducted our analysis using all three TFM-Tokenizer variants and the neural tokenizer from LaBraM (Jiang et al., 2024), testing them on the test splits of both the TUEV and IIIC datasets, which have multiple classes. All tokenizers employed a fixed vocabulary size of $8,192$ tokens for consistency and fair comparison.
[…]
Figure 3a visualizes the class-token uniqueness scores for each class in both datasets. A robust tokenizer should capture class-distinctive tokens across all dataset classes through unsupervised pretraining. To assess this, we computed the geometric mean (GM) of class-token uniqueness scores, as shown in Table 2. Our TFM-Tokenizer reduces token utilization by more than two-fold compared to the neural tokenizer on TUEV ($21.13\%\rightarrow 9.78\%$) and nearly two-fold on IIIC ($15.25\%\rightarrow 8.26\%$). It also significantly improves learning of class-unique tokens compared to neural tokenizer ($0.034\%\rightarrow 2.14\%$on TUEV, $0.0\%\rightarrow 1.429\%$ on IIIC). These results demonstrate that the TFM-Tokenizer captures more compact and useful tokens than the neural tokenizer. Additionally, TFM-Tokenizer achieves a higher class-token uniqueness score across all classes compared to TFM-Tokenizer-R ($0.0\%\rightarrow 1.429\%$ on IIIC) and TFM-Tokenizer-S ($0.619\%\rightarrow 1.429\%$ on IIIC), as depicted in Figure 3a. This further validates joint frequency-temporal modeling in EEG analysis.
Q2.2 - Tokens for Similar-Class Sample Mining: We conducted an EEG signal mining experiment based on similar-class sample retrieval. Given a multi-channel EEG sample, we first obtain its discrete token representation. Using the Jaccard similarity score, we then retrieve the top $K$ most similar samples from the dataset and compute the precision score for correctly retrieving samples of the same class. For this study, we constructed a balanced subset from the IIIC and TUEV datasets and tested all four tokenization methods. The retrieval performance, illustrated in Figure 3b, shows that all TFM-Tokenizer variants significantly outperform neural tokenizer. Notably, TFM-Tokenizer-S and TFM-Tokenizer achieve nearly $60\%$ precision on the TUEV for $K=1$. While the Jaccard similarity measure demonstrates initial feasibility, further research is needed to identify optimal metrics for token-based EEG retrieval.
Figure 3. Analysis of token quality across three TFM-Tokenizer variants and the neural tokenizer. (a) Comparison of class-token uniqueness scores across all classes. (b) Retrieval performance comparison of tokenizers in a similar-class sample mining task.
Q2.3 - Evaluating the Frequency Learning of TFM-Tokenizer Tokens: In this experiment, we compare the frequency and temporal-domain encoders of the TFM-Tokenizer to evaluate their ability to capture diverse frequency features in EEG signals. Specifically, we arrange all tokens in temporal order and perform a discrete Fourier transform on the token sequence. This process decomposes the tokens into frequencies, where each frequency reflects the degree of change between tokens at various scales. Larger changes indicate more diverse token representations. Then, we compute spectral entropy, defined as the normalized Shannon entropy of the amplitude values, to quantify how energy is distributed across the spectrum. Higher spectral entropy means that the model has learned a broader range of frequency features, capturing differences from both large-scale trends and fine details. Figure 4 shows that on the TUEV, TUAB, and CHBMIT datasets, the frequency encoder produces tokens with significantly higher spectral entropy than the temporal encoder. For example, on the TUEV dataset, the frequency encoder achieved an average spectral entropy of 0.26, while the temporal encoder reached only 0.14. This multi-scale sensitivity benefits downstream tasks such as classification, where learning detailed differences in EEG tokens can improve performance.
@@
Figure 4. An analysis of how the proposed frequency and temporal-domain encoders capture frequency features, by using the spectral entropy of the learned token sequences from randomly selected samples. Higher values indicate that the tokens contain richer frequency information.
5.4. Q3: Does TFM-Tokenizer Enhance LaBraM?
Table 3. Performance Comparison of LaBraM with their neural tokenizer vs TFM-Tokenizer
Dataset Tokenizer Performance Metrics
Balanced Acc. Cohen’s Kappa Weighted F1
TUEV Neural Tokenizer $0.4682\pm 0.0856$ $0.5067\pm 0.0413$ $0.7466\pm 0.0202$
TFM-Tokenizer 0.5147±$\pm 0.0174$uparrow 0.5220±$\pm 0.0153$uparrow 0.7533±$\pm 0.0094$uparrow
@@ 5.4. Q3: Does TFM-Tokenizer Enhance LaBraM?
TUAB Neural Tokenizer $0.7720\pm 0.0046$ $0.8498\pm 0.0036$ $0.8534\pm 0.0027$
TFM-Tokenizer 0.7765±$\pm 0.0016$uparrow 0.8518±$\pm 0.0051$uparrow 0.8584±$\pm 0.0022$uparrow
To assess the scalability of TFM-Tokenizer, we investigated its ability to enhance an existing EEG foundation model. We selected LaBraM (Jiang et al., 2024), which employs a neural tokenizer solely for pretraining. This setup makes it an ideal candidate for this study. We replaced LaBraM neural tokenizer with TFM-Tokenizer during the masked EEG modeling stage and evaluated its performance on TUEV and TUAB, presented in Table 3. On TUEV, LaBraM with TFM-Tokenizer achieves a $9\%$ increase in balanced accuracy ($0.4682\rightarrow 0.5147$) and a $3\%$ increase in Cohen’s Kappa ($0.5067\rightarrow 0.5220$). On TUAB, TFM-Tokenizer consistently outperforms the neural tokenizer. These results confirm the capability of TFM-Token in enhancing the performance of EEG foundation models. The increase in balanced accuracy suggests that our tokenizer learns more class-discriminative tokens than the neural tokenizer.
Figure 5. Overview of motifs captured by TFM-Tokenizer on the TUEV dataset: (a) presents three samples from the PLED class, while (b) displays three samples from the GPED class.
5.5. Q4: Interpretability of Learned Tokens
We conducted a visual inspection study to determine whether the learned tokens by our TFM-Tokenizer capture and represent distinct time-frequency motifs in EEG signals. Our findings are presented in Figure 5, which illustrates the unique tokens identified by TFM-Tokenizer on the TUEV dataset. Each token represents a spectral window and its corresponding raw EEG patch (1s window with 0.5s overlap). For more precise visualization, we highlight each class’s most frequently occurring tokens, assigning different colors to each token. This analysis confirms that our TFM-Tokenizer effectively captures class-specific distinct EEG patterns, encoding them into discrete tokens. For instance, token $4035$ in the PLED class consistently captures a characteristic drop followed by a rise in EEG signals, maintaining its structure across different samples despite variations in noise, amplitude, and minor shifts within the window. Similarly, tokens such as $5096$ and $3751$ in the GPED class demonstrate the advantages of combining frequency and temporal features, as they remain robust to minor temporal shifts and warping within a window due to emphasizing spectral patterns. However, we also identified certain limitations associated with using fixed windowing and overlapping segments for tokenization. Specifically, when large shifts cause a distinct EEG pattern to be split across two adjacent windows, the model may assign them separate tokens, potentially treating them as distinct patterns.
6. Conclusion
We introduced TFM-Token, a fully discrete tokenization framework consisting of TFM-Tokenizer and TFM-Encoder modules. TFM-Tokenizer effectively encodes time-frequency motifs from single-channel EEG into distinct tokens, which TFM-Encoder utilizes for downstream tasks. Comprehensive evaluations across multiple datasets demonstrate that TFM-Token outperforms state-of-the-art models while using fewer parameters. Our analysis confirms the robustness, class distinctiveness, and interpretability of learned tokens, providing deeper insights into EEG tokenization. These findings lay a strong foundation for future research in developing tokens that extend beyond fixed window segments and enhance capturing temporal-frequency information.
Appendix
[…]
1Contents
Appendix A Additional Results
A.1. Performance on CHB-MIT and IIIC Seizure
Table 4 and 5 presents the performance comparison of TFM-Token with baselines on seizure detection (CHB-MIT) and seizure type classification (IIIC Seizure) tasks. TFM-Token outperforms all baselines across all metrics in both datasets. On the CHB-MIT dataset with a highly imbalanced binary classification task, BIOT is the only baseline with an AUC-PR above $0.25$. However, TFM-Token surpasses BIOT, achieving an $8\%$ improvement in AUC-PR ($0.3127\rightarrow 0.3379$) and a $4.5\%$ increase in AUROC ($0.8456\rightarrow 0.8839$), demonstrating better robustness to class imbalance. For the IIIC Seizure dataset, where the task is to classify 10-second, 16-channel EEG segments into six classes, TFM-Token improves Cohen’s Kappa by $9.5\%$ ($0.4549\rightarrow 0.4985$) and Weighted F1 by $8.5\%$ ($0.5387\rightarrow 0.5847$) over ContraWR, which achieves second best results.
The superior performance of TFM-Token across four EEG datasets shows the promise of a fully discretized framework that has the potential to enhance future EEG foundation models. These results also underscore the importance of capturing both temporal and frequency information, highlighting the critical role of frequency learning in EEG analysis.
Table 4. Seizure detection performance comparison on the CHB-MIT dataset
Models Number CHB-MIT (seizure detection)
of Params Balanced Acc. AUC-PR AUROC
SPaRCNet(Jing et al., 2023) 0.79M $0.5483\pm 0.0164$ $0.1264\pm 0.0112$ $0.7953\pm 0.0402$
ContraWR(Yang et al., 2023) 1.6M $0.6479\pm 0.0423$ $0.2312\pm 0.0316$ $0.8198\pm 0.0197$
CNN-Transformer(Peh et al., 2022) 3.2M $0.5900\pm 0.0429$ $0.2250\pm 0.0512$ $0.8333\pm 0.0326$
FFCL(Li et al., 2022) 2.4M $0.6565\pm 0.0260$ $0.2115\pm 0.0286$ $0.8071\pm 0.0117$
ST-Transformer(Song et al., 2021) 3.5M $0.6717\pm 0.0271$ $0.2133\pm 0.0148$ $0.8657\pm 0.0139$
BIOT(Yang et al., 2024) 3.2M $0.6582\pm 0.0896$ $0.3127\pm 0.0890$ $0.8456\pm 0.0333$
LaBraM-Base(Jiang et al., 2024) 5.8M $0.5035\pm 0.0078$ $0.0959\pm 0.0742$ $0.6624\pm 0.1050$
TFM-Token 1.9M 0.6750 $\pm$ 0.0392 0.3379 $\pm$ 0.0515 0.8839 $\pm$ 0.0173
Table 5. Seizure type classification performance comparison on the IIIC Seizure dataset
Models Number IIIC Seizure (seizure type classification)
of Params Balanced Acc. Cohen’s Kappa Weighted F1
SPaRCNet(Jing et al., 2023) 0.79M $0.5011\pm 0.0286$ $0.4115\pm 0.0297$ $0.4996\pm 0.0262$
ContraWR(Yang et al., 2023) 1.6M $0.5421\pm 0.0123$ $0.4549\pm 0.0166$ $0.5387\pm 0.0138$
CNN-Transformer(Peh et al., 2022) 3.2M $0.5395\pm 0.0144$ $0.4500\pm 0.0165$ $0.5413\pm 0.0176$
FFCL(Li et al., 2022) 2.4M $0.5309\pm 0.0217$ $0.4412\pm 0.0253$ $0.5315\pm 0.0277$
ST-Transformer(Song et al., 2021) 3.5M $0.5093\pm 0.0122$ $0.4217\pm 0.0151$ $0.5217\pm 0.0110$
BIOT(Yang et al., 2024) 3.2M $0.4458\pm 0.0183$ $0.3418\pm 0.0228$ $0.4511\pm 0.0207$
LaBraM-Base(Jiang et al., 2024) 5.8M $0.4736\pm 0.0101$ $0.3716\pm 0.0128$ $0.4765\pm 0.0097$
TFM-Token 1.9M 0.5775 $\pm$ 0.0042 0.4985 $\pm$ 0.0039 0.5847 $\pm$ 0.0050
A.2. Effect of Masked Token Prediction in EEG Tokenization
@@
We conducted an ablation study on TFM-Encoder to assess the impact of masked token prediction pretraining in a fully discretized framework. Using a pretrained TFM-Tokenizer, we compared two approaches: (1) masked token prediction pretraining followed by fine-tuning and (2) direct fine-tuning without pretraining. This experiment was performed on the TUEV dataset across all three TFM-Tokenizer variants, with results summarized in Figure 6. While Cohen’s Kappa and Weighted F1 showed no significant differences between the two approaches, masked token prediction pretraining significantly improved balanced accuracy across all TFM-Tokenizer variants. This suggests that pretraining enhances class-wise prediction consistency by capturing token dependencies and making TFM-Encoder more robust to missing channels or time segments, a common challenge in EEG analysis.
Figure 6. Masked Token Prediction Ablation
A.3. Removing Position Embedding in TFM-Tokenizer Improves Token Learning
@@
Table 6. TFM-Tokenizer Comparison with and w/o Position Embedding (PE) on TUEV Dataset
@@
TFM-Tokenizer w/o PE $9.78$ $2.14$ 0.4943 ±$\pm 0.052$ 0.5337 ±$\pm 0.031$ 0.7570 ±$\pm 0.016$
Through our empirical analysis, we found that the performance of our TFM-Token significantly improved when no position embedding was applied to the TFM-Tokenizer. EEG patterns are inherently chaotic and non-stationary, meaning similar motifs can occur at any position within the signal. An ideal tokenizer should be capable of capturing and representing such EEG motifs as distinct tokens without relying on positional information.
@@
A.4. Ablation on Token Vocabulary Size

### v2

@@ C.1 Additional Results on Token Quality Analysis and Frequency Learning
Figure 7: An analysis of how the proposed frequency and temporal-domain encoders capture frequency features, by using the spectral entropy of the learned token sequences from randomly selected samples. Higher values indicate that the tokens contain richer frequency information.
In this section, we present more results on token quality analysis, specifically focusing on token utilization and frequency learning capability of our tokenizer. Additional token uniqueness and consistency experiments on IIIC dataset is presented in Figure 6b and c.
Token utilization: Token utilization ($\%$) score was calculated as the percentage of unique tokens activated from the total available vocabulary size. Additionally, we computed the geometric mean (GM) of class-token uniqueness scores along with the utilization score, and the results are presented in Table 6. Our TFM-Tokenizer reduces token utilization by more than two-fold compared to the neural tokenizer on TUEV ($21.13\%\rightarrow 9.78\%$) and nearly two-fold on IIIC ($15.25\%\rightarrow 8.26\%$). It also significantly improves learning of class-unique tokens compared to the neural tokenizer ($0.034\%\rightarrow 2.14\%$ on TUEV, $0.0\%\rightarrow 1.429\%$ on IIIC).
Evaluating the Frequency Learning of TFM-Tokenizer Tokens: In this experiment, we compare the frequency and temporal-domain encoders of the TFM-Tokenizer to evaluate their ability to capture diverse frequency features in EEG signals. Specifically, we arrange all tokens in temporal order and perform a discrete Fourier transform on the token sequence. This process decomposes the tokens into frequencies, where each frequency reflects the degree of change between tokens at various scales. Larger changes indicate more diverse token representations. Then, we compute spectral entropy, defined as the normalized Shannon entropy of the amplitude values, to quantify how energy is distributed across the spectrum. Higher spectral entropy means that the model has learned a broader range of frequency features, capturing differences from both large-scale trends and fine details. Figure 7 shows that on the TUEV, TUAB, and CHBMIT datasets, the frequency encoder produces tokens with significantly higher spectral entropy than the temporal encoder. For example, on the TUEV dataset, the frequency encoder achieved an average spectral entropy of 0.26, while the temporal encoder reached only 0.14. This multi-scale sensitivity benefits downstream tasks such as classification, where learning detailed differences in EEG tokens can improve performance.
C.2 Additional results on Frequency and Temporal Modeling for EEG Tokenization
Table 7: Ablation study on input representation to TFM-Tokenizer
Models TUEV (event type classification) TUAB (abnormal detection)
Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC
TFM-Tokenizer-R $\underline{0.4898}\pm 0.0105$ $0.5194\pm 0.0195$ $0.7518\pm 0.0095$ $\underline{0.8033}\pm 0.0021$ $\underline{0.8908}\pm 0.0027$ $\underline{0.8849}\pm 0.0024$
TFM-Tokenizer-S $0.4708\pm 0.0339$ $\underline{0.5275}\pm 0.0314$ $\underline{0.7538}\pm 0.0152$ $0.7927\pm 0.0044$ $0.8814\pm 0.0095$ $0.8836\pm 0.0052$
TFM-Tokenizer 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163 0.8152 $\pm$ 0.0014 0.8946 $\pm$ 0.0008 0.8897 $\pm$ 0.0008
1. The best results are bolded, while the second-best are underlined.
In Table 7 we provide detailed results of our ablation study discussed under Section 4.5.
C.3 Token Generalization Assessment through Cross-Dataset Experiments
Table 8: Cross dataset generalizability experiments under single dataset settings
Testing Tokenizer MTP Performance Metrics
Dataset Dataset Dataset Balanced Acc. Cohen’s Kappa Weighted F1
TUEV TUEV TUEV 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163
IIIC TUEV 0.4722 $\pm$ 0.0578 0.4990 $\pm$ 0.0237 0.7380 $\pm$ 0.0137
IIIC 0.4291 $\pm$ 0.0235 0.5195 $\pm$ 0.0200 0.7534 $\pm$ 0.0100
TUAB TUEV 0.4651 $\pm$ 0.0449 0.5925 $\pm$ 0.0249 0.7847 $\pm$ 0.0136
TUAB 0.5252 $\pm$ 0.0431 0.6187 $\pm$ 0.0285 0.8018 $\pm$ 0.0138
CHB-MIT TUEV 0.4979 $\pm$ 0.0444 0.5995 $\pm$ 0.0225 0.7885 $\pm$ 0.0122
CHB-MIT 0.5898 $\pm$ 0.0192 0.6591 $\pm$ 0.0106 0.8196 $\pm$ 0.0045
To evaluate the robustness of our tokenizer, we conducted cross-dataset experiments under two settings: (1) fixing the tokenizer and performing masked token prediction (MTP) & finetuning on a different target dataset and (2) fixing the tokenizer and MTP, followed by finetuning TFM-Encoder only on the target dataset. Results are presented in Table 8, which demonstrates strong generalizability, with our TFM-Tokenizer achieving the best performance on TUEV when pretrained on CHBMIT—outperforming the best-reported result in four dataset settings. These findings highlight the potential of our tokenizer as a foundation for a scalable, universal EEG tokenizer.
C.4 Additional Results on TFM-Tokenizer Improving Existing Foundation Models
Table 9: Performance comparison of LaBraM and BIOT with and w/o our TFM-Tokenizer.
Dataset Exp. Method Performance Metrics
Setting Balanced Acc. Cohen’s Kappa Weighted F1
TUEV Single BIOT 0.4679 $\pm$ 0.0354 0.4890 $\pm$ 0.0407 0.7352 $\pm$ 0.0236
BIOT-TFM 0.4228 $\pm$ 0.0162 0.5230 $\pm$ 0.0226 $\uparrow$ 0.7490 $\pm$ 0.0114 $\uparrow$
LaBraM 0.4682 $\pm$ 0.0856 0.5067 $\pm$ 0.0413 0.7466 $\pm$ 0.0202
LaBraM-TFM 0.5147 $\pm$ 0.0174 $\uparrow$ 0.5220 $\pm$ 0.0153 $\uparrow$ 0.7533 $\pm$ 0.0094 $\uparrow$
Multiple BIOT 0.5281 $\pm$ 0.0225 0.5273 $\pm$ 0.0249 0.7492 $\pm$ 0.0082
BIOT-TFM 0.5530 $\pm$ 0.0089 $\uparrow$ 0.5464 $\pm$ 0.0137 $\uparrow$ 0.7625 $\pm$ 0.0069 $\uparrow$
LaBraM 0.5550 $\pm$ 0.0403 0.5175 $\pm$ 0.0339 0.7450 $\pm$ 0.0194
LaBraM-TFM 0.5541 $\pm$ 0.0316 0.5367 $\pm$ 0.0281 $\uparrow$ 0.7567 $\pm$ 0.0165 $\uparrow$
IIIC Single BIOT 0.4458 $\pm$ 0.0183 0.3418 $\pm$ 0.0228 0.4511 $\pm$ 0.0207
BIOT-TFM 0.4633 $\pm$ 0.0078 $\uparrow$ 0.3663 $\pm$ 0.0103 $\uparrow$ 0.4689 $\pm$ 0.0090 $\uparrow$
LaBraM 0.4736 $\pm$ 0.0101 0.3716 $\pm$ 0.0128 0.4765 $\pm$ 0.0097
LaBraM-TFM 0.4814 $\pm$ 0.0075 $\uparrow$ 0.3795 $\pm$ 0.0091 $\uparrow$ 0.4841 $\pm$ 0.0062 $\uparrow$
Multiple BIOT 0.4414 $\pm$ 0.0035 0.3362 $\pm$ 0.0040 0.4483 $\pm$ 0.0033
BIOT-TFM 0.5050 $\pm$ 0.0037 $\uparrow$ 0.4098 $\pm$ 0.0052 $\uparrow$ 0.5139 $\pm$ 0.0025 $\uparrow$
LaBraM 0.4736 $\pm$ 0.0037 0.3658 $\pm$ 0.0033 0.4708 $\pm$ 0.0015
LaBraM-TFM 0.4782 $\pm$ 0.0065 $\uparrow$ 0.3737 $\pm$ 0.0076 $\uparrow$ 0.4790 $\pm$ 0.0082 $\uparrow$
@@ C.4 Additional Results on TFM-Tokenizer Improving Existing Foundation Models
CHB-MIT Single BIOT 0.6582 $\pm$ 0.0896 0.3127 $\pm$ 0.0890 0.8456 $\pm$ 0.0333
BIOT-TFM 0.5893 $\pm$ 0.0197 0.3341 $\pm$ 0.0349 $\uparrow$ 0.8752 $\pm$ 0.0123 $\uparrow$
LaBraM 0.5035 $\pm$ 0.0078 0.0959 $\pm$ 0.0742 0.6624 $\pm$ 0.1050
LaBraM-TFM 0.5473 $\pm$ 0.047 $\uparrow$ 0.2376 $\pm$ 0.0461 $\uparrow$ 0.7863 $\pm$ 0.0438 $\uparrow$
Multiple BIOT 0.7068 $\pm$ 0.0457 0.3277 $\pm$ 0.0460 0.8761 $\pm$ 0.0284
BIOT-TFM 0.6197 $\pm$ 0.0085 0.3484 $\pm$ 0.0078 $\uparrow$ 0.8726 $\pm$ 0.0098
LaBraM 0.5260 $\pm$ 0.0369 0.2138 $\pm$ 0.0523 0.7750 $\pm$ 0.0540
LaBraM-TFM 0.5579 $\pm$ 0.0394 $\uparrow$ 0.2445 $\pm$ 0.0351 $\uparrow$ 0.7887 $\pm$ 0.0423 $\uparrow$
Table 9 presents detailed results on integrating TFM-Tokenizer with BIOT and LaBraM. Across all metrics and settings, TFM-Tokenizer improves performance in $93\%$ of cases, demonstrating its effectiveness in enhancing existing EEG foundation models.
C.5 Effect of Masked Token Prediction in EEG Tokenization
Figure 8: Masked Token Prediction Ablation
@@ C.5 Effect of Masked Token Prediction in EEG Tokenization
We conducted an ablation study on downstream transformer to assess the impact of masked token prediction pretraining in a fully discretized framework. Using a pretrained TFM-Tokenizer, we compared two approaches: (1) masked token prediction pretraining followed by fine-tuning and (2) direct fine-tuning without pretraining. This experiment was performed on the TUEV dataset across all three TFM-Tokenizer variants, with results summarized in Figure 8. While Cohen’s Kappa and Weighted F1 showed no significant differences between the two approaches, masked token prediction pretraining significantly improved balanced accuracy across all TFM-Tokenizer variants. This suggests that pretraining enhances class-wise prediction consistency by capturing token dependencies and making downstream transformer more robust to missing channels or time segments, a common challenge in EEG analysis.
C.6 Removing Position Embedding in TFM-Tokenizer Improves Token Learning
@@
Table 10: TFM-Tokenizer Comparison with and w/o Position Embedding (PE) on TUEV Dataset
@@
TFM-Tokenizer w/o PE $9.78$ $2.14$ 0.4943 $\pm 0.052$ 0.5337 $\pm 0.031$ 0.7570 $\pm 0.016$
Through our empirical analysis, we found that the performance significantly improved when no position embedding was applied to the TFM-Tokenizer. EEG patterns are inherently chaotic and non-stationary, meaning similar motifs can occur at any position within the signal. An ideal tokenizer should be capable of capturing and representing such EEG motifs as distinct tokens without relying on positional information.
@@
C.7 Downstream Model Ablation
Table 11: Ablation on number of transformer layers in the downstream model
# Number of Performance Metrics
Layers Params. Balanced Acc. Cohen’s Kappa Weighted F1
1 0.58M 0.4486 $\pm$ 0.0297 0.5404 $\pm$ 0.0168 0.7603 $\pm$ 0.0096
2 0.63M 0.4920 $\pm$ 0.0595 0.5758 $\pm$ 0.0169 0.7780 $\pm$ 0.0089
4 0.72M 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163
6 0.82M 0.5025 $\pm$ 0.0592 0.4996 $\pm$ 0.0208 0.7410 $\pm$ 0.0104
12 1.12M 0.5016 $\pm$ 0.0730 0.5088 $\pm$ 0.0272 0.7456 $\pm$ 0.0139
We ablated the number of transformer layers in the downstream model on the TUEV dataset, with results presented in Table 11. Notably, even with significantly fewer parameters (two layers), the model maintains competitive and, in some cases, better performance across key metrics. This highlights the potential for developing lightweight and efficient models for EEG analysis without substantial performance trade-offs.
C.8 Ablation on Token Vocabulary Size